# Autoregressive Processes on Stiefel and Grassmann Manifolds

### v1

@@ A Basic Riemannian Geometry
The geodesic curve going from $p$ to $q$ is any curve $\tau:[a,b]\rightarrow M$ such that $\operatorname{dist}_{\tau}(p,q)\leq\operatorname{dist}_{\gamma}(p,q)$ for all curves $\gamma$ starting at $p$ and ending at $q$. Moreover, the distance $\operatorname{dist}(p,q)$ between two points $p,q\in M$ is defined as the distance along the geodesic going from $p$ to $q$. Observe here that the distance function is a topological metric on $M$. Note, that $\tau$ is only guaranteed to be unique when $p$ and $q$ lie in a neighbourhood small enough. Since geodesics are locally unique one can define a chart which is locally an isometry between $T_{p}M$, with metric from the inner product $g_{p}$ and the distance function on $M$.
@@
Let $p\in M$ be a point and let $U\subseteq T_{p}M$ be open and such that all length minimizing geodesics $0,\norm{V}]\rightarrow M$ with initial velocity $X=V/\norm{V}$, $V\in U$ are unique. Then, the assignment $\exp_{p}(V)=\gamma_{p}(V)$ where $\gamma_{p}$ is the unique geodesic starting at $p$ with initial velocity $V/\norm{V}$ and length $\norm{V}$, is a map $U\rightarrow M$ called the Riemannian exponential map at $p$. It’s inverse $\log_{p}:\exp_{p}(U)\rightarrow U$ is the geodesic normal chart around $p$, and it’s associated coordinates are called the geodesic normal coordinates.
The geodesic normal coordinates are of particular importance. In geodesic normal coordinates, the gradient and Hessian are reduced to the Euclidean gradient and Hessian. That is, if $x_{i}$ are geodesic normal coordinates centred at $p$, i.e. the functions generated by the exponential map and some orthonormal basis on $T_{p}M$, then
@@
Utilizing spherical geodesic normal coordinates centred around $p$,
@@ B Simulations
For the case when $M=O(n)$ we apply the algorithm as outlined in Section 4.1, i.e. through the barycentre algorithm compute the average of all the $Z_{j+1}^{-1}Z_{j}$. The error between the estimate and $\Phi$ is of the order $\mathcal{O}(\sqrt{N}^{-1})$ with coefficient depending on all other parameters, $n,\Phi,\sigma$. This is expected since this algorithm is a generalization of sample average estimates in Euclidean space, which also has the same order term. In figure 1 this is illustrated for the case $O(20)$ and $\sigma=0.1$.
Figure 1: Simulation an AR(1) process using barycentres on $O_{20}$ where the noise is of approximate amplitude $10^{-1}$.
@@
The opposite behaviour can be seen when increasing the parameter $k$ in $\operatorname{St}_{n,k}$ when $k\lesssim n/2$ as seen in Figure 5 for $\operatorname{St}_{20,k}$ and $N=200$. This is the case even though the dimension of $\operatorname{St}_{n,k}$ increases, the degrees of freedom of the parameter $\Phi$ decreases when $k$ increases.
@@
For the same reason as for the Stiefel manifolds, higher degrees of freedom for $\Phi$ means larger error. Fixing $k$ we see the same behaviour as previously, see Figure 8 for the case $k=5$.
@@
In Figure 9 we observe the same behaviour as in the case of the Stiefel manifold, for the same reason.

### v2

@@ Appendix A Basic Riemannian Geometry
The geodesic curve going from $p$ to $q$ is any curve $\tau:[a,b]\rightarrow M$ such that $\operatorname{dist}_{\tau}(p,q)\leq\operatorname{dist}_{\gamma}(p,q)$ for all curves $\gamma$ starting at $p$ and ending at $q$. Moreover, the distance $\operatorname{dist}(p,q)$ between two points $p,q\in M$ is defined as the distance along the geodesic going from $p$ to $q$. Observe here that the distance function is a topological metric on $M$. Note, that $\tau$ is only guaranteed to be unique when $p$ and $q$ lie in a neighborhood small enough. Since geodesics are locally unique one can define a chart which is locally an isometry between $T_{p}M$, with metric from the inner product $g_{p}$ and the distance function on $M$.
@@
Let $p\in M$ be a point and let $U\subseteq T_{p}M$ be open and such that all length minimizing geodesics $0,\norm{V}]\rightarrow M$ with initial velocity $X=V/\norm{V}$, $V\in U$ are unique. Then, the assignment $\exp_{p}(V)=\gamma_{p}(V)$ where $\gamma_{p}$ is the unique geodesic starting at $p$ with initial velocity $V/\norm{V}$ and length $\norm{V}$, is a map $U\rightarrow M$ called the Riemannian exponential map at $p$. It’s inverse $\log_{p}:\exp_{p}(U)\rightarrow U$ is the geodesic normal chart around $p$, and it’s associated coordinates are called the geodesic normal coordinates. The maximal $\rho_{p}>0$ such that $\exp_{p}$ is injective when it is restricted to the ball
$B_{\rho}(0)=\{x\in T_{p}X:\norm{x}<\rho\}$
is referred to as the injectivity radius of $M$ at $p$. The minimal injectivity radius at all points on $M$, $\rho=\inf_{p\in M}\rho_{p}$ is referred to as the injectivity radius of $M$.
The geodesic normal coordinates are of particular importance. In geodesic normal coordinates, the gradient and Hessian are reduced to the Euclidean gradient and Hessian. That is, if $x_{i}$ are geodesic normal coordinates centered at $p$, i.e. the functions generated by the exponential map and some orthonormal basis on $T_{p}M$, then
@@
Utilizing spherical geodesic normal coordinates centered around $p$,
@@ Appendix B Simulations
For the case when $M=O(n)$ we apply the algorithm as outlined in Section 4.1, i.e. through the barycenter algorithm compute the average of all the $Z_{j+1}^{-1}Z_{j}$. The error between the estimate and $\Phi$ is of the order $\mathcal{O}(\sqrt{N}^{-1})$ with coefficient depending on all other parameters, $n,\Phi,\sigma$. This is expected since this algorithm is a generalization of sample average estimates in Euclidean space, which also has the same order term. In figure 1 this is illustrated for the case $O(20)$ and $\sigma=0.1$.
Figure 1: Simulation an AR(1) process using barycenters on $O_{20}$ where the noise is of approximate amplitude $10^{-1}$.
@@
The opposite behavior can be seen when increasing the parameter $k$ in $\operatorname{St}_{n,k}$ when $k\lesssim n/2$ as seen in Figure 5 for $\operatorname{St}_{20,k}$ and $N=200$. This is the case even though the dimension of $\operatorname{St}_{n,k}$ increases, the degrees of freedom of the parameter $\Phi$ decreases when $k$ increases.
@@
For the same reason as for the Stiefel manifolds, higher degrees of freedom for $\Phi$ means larger error. Fixing $k$ we see the same behavior as previously, see Figure 8 for the case $k=5$.
@@
In Figure 9 we observe the same behavior as in the case of the Stiefel manifold, for the same reason.